A Table Affair, which sells European tablecloths at local farmers markets, has the following weekly revenue and cost functions:
R(t) = 120t
C(t) = 40t + 700
where R(t) represents the weekly revenue in dollars for selling t tablecloths and C(t) represents the weekly costs in dollars for selling t tablecloths. Find the number of tablecloths A Table Affair must sell to break-even or make a profit.
A Table Affair must sell at least______table cloths to break even or make a profit.

Answer by toddcarnes(6)

First you need to recognize that to break even means that R(t) = C(t).
Therefore, let 120t = 40t + 700 and then solve for t as follows.
120t = 40t + 700
120t - 40t = 700
80t = 700
t = 700 / 80
t = 8.75
In other words, you would need to sell 8.75 tablecloths to break even. That means you need to sell at least 9 tablecloths to make a profit.